QUES 1a: Rani and Co. manufactures automobile accessories and parts. The following are the total
processing costs for each unit.

                                                                                    $

Direct material                                                           5,000

Direct labour                                                               8,000

Variable Factory Overhead                                        6,000

Fixed factory overhead                                               50,000

The same units are available in the local market. The purchase price of the component is $ 22,000 per unit. The fixed overhead would continue to be incurred even when the component is bought from outside, although there would be reduction to the extent of $2,000 per unit. However, this reduction does not occur, if the machinery is rented out.

Required:

(i) Should the part be made or bought, considering that the present capacity when released would remain idle?

(i) In case, the released capacity can be rented out to another manufacturer for $4,500 per unit, what should be the decision?

b. Flowers Inc. manufactures silk roses. Bud Company has approached Flowers with a proposal to buy 2,000 silk roses for $4.00 each. Regular customers are charged $4.25 for each rose. Flowers has the necessary capacity. The following costs are associated annually with silk roses with the company's normal production and sales of 10,000 roses:

   $

Direct material

21,000

Direct labor

13,000

Manufacturing overhead

   9,000

Total

43,000

Forty percent of the manufacturing overhead is variable. All fixed overhead is allocated equally to all products produced.

  1. Should Flowers accept the order from Bud Company?
  2. What qualitative (non-financial) considerations may influence your decision? *Give at least 4 qualitative reasons
Answer #1

Relevant cost of making = Avoidable cost

= 5000+8000+6000+2000

= $21,000

Purchase cost = $22,000

1.Hence, the part should be made

2.Buy, since net benefit per unit = 4500-1000

= $3500

a.Relevant cost per unit = 2.1+1.3+0.9*40% = $3.76

The order should be accepted since price is higher than relevant cost

b.Effect on regular sales

potential of long term relations with the party giving special order